OpenStudy (anonymous):

-1/8c = 2 ? multiplication and divison equations

OpenStudy (anonymous):

-1/8c = 2 => -1= 16 c c = -1/16

OpenStudy (anonymous):

thanks you think you can help me on some more ? @ashsek

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

-5/8c = 20

OpenStudy (anonymous):

-5 / 8c = 20 => -5 = 160c c = -5 / 160 = -1 / 32

OpenStudy (anonymous):

3m ____ 7 = 12 m= ?

OpenStudy (anonymous):